tickets to a high scholl play cost $1.10 for each adult and 40 cents for each child. if 360 tickets are sold for a total of $282.60 how many tickets of each kind are sold?
Answer by vhardiker(2)

Suppose x Adult and y Child Tickets are sold
Thus, from the information given we know
x + y = 360 (total # of tickets sold) ----- Equation 1
1.1*x + 0.4*y = 282.6 ----- Equation 2
Thus we have two equations and two unknowns
From Equation 1 we have x + y = 360 which implies y = 360 - x
Substitute this in Equation 2 to obtain
1.1*x + 0.4*(360 - x) = 282.6
which implies 1.1*x + 0.4*360 - 0.4*x =282.6
Now keep the unknown x on the left hand side of the above equation and take the
known 0.4*360 to the right had side, we obtain
1.1*x - 0.4*x = 282.6 - 0.4*360 (sign of 0.4*360 changes from + on left hand side to - when brought to the right hand side)
=> 0.7*x = 282.6 - 144
=> 0.7*x = 138.6
=> x = 138.7/0.7 = 198
Thus x = 198 and from Equation 1 we have 198 + y = 360
=> y = 360 - 198 = 162
Answer : 198 Adult Tickets and 162 Child tickets were sold